OpenStudy (anonymous):

Using the properties of equality to solve the equation -2b + 7 = -13, you would _____. a. add 13 and then divide by -2 b.add 13 and then add 2 c.subtract 7 and then add 2 d.subtract 7 and then divide by -2

OpenStudy (anonymous):

I'm really stuck on this one

Directrix (directrix):

Let's work it and see. -2b + 7 = -13 Subtract 7 from both sides -2b + 7 = -13 -7 -7 -2b = 6 Divide by -2 b = -3 @TeenFishermen Look at this solution and then choose the correct option, okay?